# How does the successive buckling events affect a galaxy bar and stellar disk? : Potential Observable Signatures For Spotting the Buckling Action -I

### v1

@@ Abstract
Until now, observations have caught up only a handful of galaxies in ongoing buckling action. Interestingly, N-body simulations over the past several decades show that almost every bar buckles or vertically thickens as soon as it reaches its peak strength during its evolution and leads to box/peanut/x (BPX) shapes. In order to understand the effect of multiple buckling events on the observable properties of galactic bar and disk, we perform an N-body simulation of a Milky Way-type disk. The axisymmetric galaxy disk forms a bar within a Gyr of its evolution and the bar undergoes two successive buckling events. We report that the time spans of these two buckling events are 220 Myr and 1 Gyr which have almost similar strengths of the bending modes. As a result of these two buckling events, the sizes of BPX shapes are around 5.8 kpc and 8.6 kpc which are around two-thirds of the bar length at the end of each buckling event. We find that the first buckling occurs at a smaller scale (radius $<$ 3 kpc) with a shorter time span affecting the larger length scales of the disk which is quantified in terms of changes in $m=$2 and $m=$ 4 Fourier modes. While the second buckling occurs at larger scales (radius $\approx$ 6 kpc) affecting the inner disk the most. Finally, we provide observable kinematic signatures (i.e. quadrupolar patterns of the line-of-sight velocities) which can potentially differentiate the successive buckling events.
@@
pubyear: 2015pagerange: How does the successive buckling events affect a galaxy bar and stellar disk? : Potential Observable Signatures For Spotting the Buckling Action -IA
@@ 1 Introduction
Most of the disk galaxies in the local universe are barred ones which have been observed mainly in the optical and infrared wavelengths (Eskridge et al., 2000; Menéndez-Delmestre et al., 2007; Sheth et al., 2008; Erwin, 2018; Lee et al., 2019). The bars can survive longer in the disk galaxies given a significantly larger bar fraction at high redshifts (Jogee et al., 2004; Sheth et al., 2008). Recent JWST observations quantify the barred galaxies at $z>2$ (Guo et al., 2023; Le Conte et al., 2024) which suggests bars play a significant role in galaxy evolution. This agrees with the high bar fraction at large redshifts in the next-generation cosmological simulations (Rosas-Guevara et al., 2020; Zhao et al., 2020; Rosas-Guevara et al., 2022; Ansar et al., 2023a, b; Kataria & Vivek, 2024; Fragkoudi et al., 2024).
@@
Over the past several decades N-body simulations have shown that the bar grows in the disk by transferring angular momentum from the inner disk to the outer disk and the surrounding dark matter halo (Athanassoula, 2003; Sellwood, 2014). The central massive concentration has been shown to suppress the bar instability (Ostriker & Peebles, 1973; Shen & Sellwood, 2004; Kataria & Das, 2018; Saha & Elmegreen, 2018; Kataria & Das, 2019; Kataria et al., 2020) while the role of halo spin on bar suppression is debated (Collier et al., 2018; Kataria & Shen, 2022; Li et al., 2023a; Chiba & Kataria, 2024; Kataria & Shen, 2024). One of the most common features associated with the bar is the BPX shape (Combes et al., 1990) that can result from the vertical thickening of planar bar orbits. The vertical thickening can result from three processes (Sellwood & Gerhard, 2020) namely; vertical buckling (Raha et al., 1991), vertical heating of 2:1 resonance (Pfenniger & Friedli, 1991; Quillen et al., 2014) and gradual trapping of 2:1 orbits resonances. The vertical buckling of the bar is similar to the bending modes in stellar disks (Sellwood & Merritt, 1994) though underlying physics is yet to be explored (Li et al., 2023b). The buckling of the bar leads to weakening in the size and strength of the bar and results in BPX shape (Debattista et al., 2005; Martinez-Valpuesta & Shlosman, 2004). The weaker buckling has been shown to destroy the bar completely (Collier, 2020).
@@ 2 Simulation of N-body galaxy model
[…]
@@ 3 Results
The top panel of Figure 1 shows the time evolution of bar strength. The bar strength continuously increases with time during the initial growth of the bar until 1.43 Gyr. The bar strength shows an abrupt decrease until 1.45 Gyr and it starts increasing continuously again until 2.2 Gyr. The bar strength does not grow between 2.2 to 3.2 Gyr though shows a short plateau distribution in this interval. Then beyond 3.2 Gyr bar strength grows and saturates in the secular evolution of the disk. Figure 2 shows the growth of bar length with time. The bar length decreases between 1.37 to 1.54 Gyr and shows a plateau between 2.2 to 3.2 Gyr like the bar strength evolution with time. The bar length continuously increases with time in secular evolution, unlike bar strength which saturates. We show the face-on and edge-on maps of bar evolution throughout the simulations in Figure 13 and discuss in Appendix A. The second panel of Figure 1 shows the time evolution of the pattern speed of the bar after the bar has formed ($A_{2}/A_{0}>0.35$). We find that pattern speed decreases with time after the bar gains strength by transferring angular momentum to the outer disk and dark matter halo (Kataria & Shen, 2022). The pattern speed shows a sharp decrease at 1.45 Gyr coinciding with the time of the sudden decrease in bar strength. In the secular phase of bar evolution, pattern speed decreases smoothly until the end of the simulation. The third panel of Figure 1 shows the time evolution of the vertical buckling strength which shows two clear peaks at 1.45 Gyr and 2.68 Gyr followed by small oscillations in secular evolution. These two buckling events, differing in 30 % in their strength, are represented by dashed and dotted vertical lines respectively which have the above-mentioned imprints on the bar strength and pattern speed evolution. The last panel of Figure 1 shows the time evolution of $\sigma_{z}/\sigma_{r}$ which is calculated within $2*R_{d}$ region of the disk. The $\sigma_{z}/\sigma_{r}$ shows a decrease as the bar grows in the disk and results in the increased radial dispersion of the disk. Afterwards, there is an abrupt and huge increase in the $\sigma_{z}/\sigma_{r}$ value ($\approx$ 0.65 to 0.77) during the first buckling at 1.45 Gyr. Further, $\sigma_{z}/\sigma_{r}$ starts decreasing until the second buckling. In the second buckling event at 2.68 Gyr, there is a slow and relatively smaller increase in $\sigma_{z}/\sigma_{r}$ values ($\approx$ 0.68 to 0.72). In the secular phase of bar evolution, $\sigma_{z}/\sigma_{r}$ decreases and saturates at the end of the simulation.
@@ 3.1 Buckling events, corresponding BPX shape lengths and kinematic signatures
Similarly, we examine the face-on stellar density with isodensity contours evolution and the quadrupolar patterns in the mean $V_{z}$ maps evolution during the second buckling in Figures 7 and 8 respectively. We find that the semi-major axis of the bar remains the same ($\approx$ 6.2 kpc) during the whole secondary buckling event. Isodenisty contours in Figure 7 show the weak spiral modes in the outer region of the disk which are not connected to the edge of the bar ($\approx$ 10 kpc). The quadrupolar patterns of the mean $V_{z}$ show a weak signal at the beginning of the buckling which becomes strengthened during the peak of buckling and fades away at around 3.08 Gyr. The region encompassing the quadrupolar patterns increases with radius from the beginning to the end of buckling event though remains confined within the semi-major axis of bar. In comparison to the first buckling, these quadrupolar patterns during the second buckling are extended, non-symmetric, and less pronounced compared to mean $V_{z}$ of the disk background.
@@ 3.2 Time Evolution of Fourier modes ($m=$ 2 and 4) throughout the disk
We observe that two buckling events have pronounced effects on the BPX shapes of bars as seen in Figure 4. In order to understand the impact of buckling throughout the disk we plot the $m=$ 2 and 4 modes in concentric rings of width equal to 0.75 kpc throughout the disk in Figures 9 and 10 respectively. The left panel of Figure 9 shows that the first buckling does not affect the growth of $A_{2}/A_{0}$ mode in the inner disk region (radius $<$ 6 kpc) while the second buckling shows a sudden dip in $A_{2}/A_{0}$ mode in this region. The division of the inner and outer regions is motivated by the difference in the time evolution of the overall pattern of $A_{2}/A_{0}$ and the dividing radius coincides with the typical bar semi-major axis length during the second buckling event. The right panel of Figure 9 shows that two buckling events have reverse trend, for a sudden dip in the time evolution of $A_{2}/A_{0}$ mode, in the outer disk (radius $>$ 6 kpc) compared to the inner part of the disk. This is interesting to observe that the first buckling which is confined to the central region of the disk (radius $<$ 3 kpc) affects the $m=2$ at the outer region of the disk. This is correlated with the growth and decay of spiral arms at the edge of bar (bar semi-major axis $\approx$ 4 kpc) during the first buckling as seen in Figure 5. On the other hand, the second buckling which spreads across the larger regions of the disk (radius $\approx$ 5 kpc) affects the $m=$ 2 mode of the central region of the disk. This is correlated with the inner region of the bar showing peanut features during the end of this buckling event as seen in Figure 7 at 3.08 Gyr. The time span of buckling events correlates inversely with the length scale of the disk where $m=$ 2 mode shows a sudden dip during its time evolution. For example the shorter time span of the first buckling shows a dip in the $m=$ 2 mode at large scale lengths of the disk and vice-versa. These modes show particular isophotal features in the face-on density maps which depend on the sequence or time span of buckling.
@@
Figure 9: Left panel shows the time evolution of m=2 Fourier mode ($A_{2}/A_{0}$) with time in concentric rings of increasing radius until 6 kpc of disk radius. The right panel show the same m=2 mode for concentric rings at a radius ranging from 6 kpc to 15 kpc. Here vertical dashed line corresponds to the first buckling ($\approx$ 1.45 Gyr) and the vertical dotted line corresponds to the second buckling event ($\approx$ 2.68 Gyr). The dashed and dotted lines corresponds to first and second buckling respectively.
@@ 4 Discussion
Observations show that only a handful of galaxies (only 8 galaxies) have been caught in ongoing buckling action (Erwin & Debattista, 2016; Li et al., 2017; Xiang et al., 2021) . Erwin & Debattista (2016) report two ongoing buckling events in their sample of 44 galaxies using photometric and kinematic signatures which suggest the buckling fraction to be around 4.5 %. Their toy model, which correctly predicts the frequency of buckling events with their observations, assumes the buckling timescale to be wide-ranged between 0.5 to 1 Gyr. Further, this study is limited by the sample size for a robust accounting of the frequency of ongoing buckling and bound on the timescale of buckling events. In a recent advancement that involves a novel approach, Xiang et al. (2021) report 5 ongoing buckling in a sample of 434 galaxies by detecting quadrupolar patterns in Los velocities in the MaNGA survey having integral field spectroscopic data (IFU). This study also conveys the message of missing detection of buckling events in their sample due to limits on signal-to-noise ratios. Their toy model, based on a semi-analytical bar evolution model, predicts a lower bound on the buckling timescale to be 130 Myr. These semi-analytical calculations do not account for the secondary buckling events which can affect the detection probability and hence the lower bound on the buckling timescales. In our N-body model, the bar undergoes two buckling episodes lasting for around 220 Myr and 1 Gyr over a time span of 9 Gyr of bar evolution. Therefore, the detection probability of buckling events must be higher than the observed events, given the bar can last for buckling around an order of magnitude larger than predicted by Xiang et al. (2021).
@@
The time span of buckling and a number of successive buckling events are the key parameters for spotting the ongoing buckling action of bars in observed galaxies. In this work, we have looked into the role of successive buckling events on the potential observable properties of bar and disk. The kinematic observable signature involves the presence of quadrupolar patterns on the Los velocity in the near face-on projection of disk. The longevity of these signatures is directly proportional to the time of buckling as evident from Figures 6 and 8. Further, the shorter time span of buckling affects $m=$ 2 and $m=$ 4 Fourier modes at larger scale lengths of disk which can provide additional signatures on the time spans of the buckling or successive buckling events in the photometric observations.
@@
Collier (2020) discussed the role of violent buckling compared to non-violent one where violent buckling has a lower timescale compared to non-violent one. The violent buckling does not destroy the bar and washes out quickly without affecting the bar in secular evolution while the non-violent one destroys the bar by sustaining bending modes for longer times. The current work has only violent buckling events ($A_{\{Buckle}>0.05$) given that do not destroy the bar irrespective of the time span of the bending mode. This highlights that the buckling mode lasting longer in the bar is not sufficient to destroy the bar.

### v2

# How do the successive buckling events affect a galaxy bar and stellar disk? : Potential Observable Signatures For Spotting the Buckling Action -I

@@ Abstract
Until now, observations have caught up only a handful of galaxies in ongoing buckling action. Interestingly, N-body simulations over the past several decades show that almost every bar buckles or vertically thickens as soon as it reaches its peak strength during its evolution and leads to box/peanut/x (BPX) shapes. In order to understand the effect of multiple buckling events on the observable properties of galactic bar and disk, we perform an N-body simulation of a Milky Way-type disk. The axisymmetric galaxy disk forms a bar within a Gyr of its evolution and the bar undergoes two successive buckling events. We report that the time spans of these two buckling events are 220 Myr and 1 Gyr which have almost similar strengths of the bending modes. As a result of these two buckling events, the full lengths of BPX shapes are around 5.8 kpc and 8.6 kpc which are around two-thirds of the full bar length at the end of each buckling event. We find that the first buckling occurs at a smaller scale (radius $<$ 3 kpc) with a shorter time span affecting the larger length scales of the disk which is quantified in terms of changes in $m=$2 and $m=$ 4 Fourier modes. While the second buckling occurs at larger scales (radius $\approx$ 6 kpc) affecting the inner disk the most. Finally, we provide observable kinematic signatures (i.e. quadrupolar patterns of the line-of-sight velocities) which can potentially differentiate the successive buckling events.
@@
pubyear: 2015pagerange: How do the successive buckling events affect a galaxy bar and stellar disk? : Potential Observable Signatures For Spotting the Buckling Action -IA
@@ 1 Introduction
Most of the disk galaxies in the local universe are barred ones which have been observed mainly in the optical and infrared wavelengths (Eskridge et al., 2000; Menéndez-Delmestre et al., 2007; Sheth et al., 2008; Erwin, 2018; Lee et al., 2019). The evidence for significant bar fractions at higher redshifts (Jogee et al., 2004; Sheth et al., 2008) suggests bars can be long-lived. Recent JWST observations quantify the barred galaxies at $z>2$ (Guo et al., 2023; Le Conte et al., 2024) which suggests bars play a significant role in galaxy evolution. This agrees with the high bar fraction at large redshifts in the next-generation cosmological simulations (Rosas-Guevara et al., 2020; Zhao et al., 2020; Rosas-Guevara et al., 2022; Ansar et al., 2023a, b; Kataria & Vivek, 2024; Fragkoudi et al., 2024).
@@
Over the past several decades N-body simulations have shown that the bar grows in the disk by transferring angular momentum from the inner disk to the outer disk and the surrounding dark matter halo (Athanassoula, 2003; Sellwood, 2014). The central massive concentration has been shown to suppress the bar instability (Ostriker & Peebles, 1973; Shen & Sellwood, 2004; Kataria & Das, 2018; Saha & Elmegreen, 2018; Kataria & Das, 2019; Kataria et al., 2020) while the role of halo spin on bar suppression is debated (Collier et al., 2018; Kataria & Shen, 2022; Li et al., 2023a; Chiba & Kataria, 2024; Kataria & Shen, 2024). One of the most common features associated with the bar is the BPX shape (Combes et al., 1990) that can result from the vertical thickening of planar bar orbits. The vertical thickening can result from three processes (Sellwood & Gerhard, 2020) namely; vertical buckling (Raha et al., 1991), vertical heating of 2:1 resonance (Pfenniger & Friedli, 1991; Quillen et al., 2014) and gradual trapping of 2:1 orbits resonances. The vertical buckling of the bar is similar to the bending modes in stellar disks (Sellwood & Merritt, 1994) though underlying physics is yet to be explored (Li et al., 2023b). The buckling of the bar leads to weakening in the size and strength of the bar and results in BPX shape (Debattista et al., 2005; Martinez-Valpuesta & Shlosman, 2004). The weaker buckling has been shown to destroy the bar completely (Collier, 2020). Recent study (Tahmasebzadeh et al., 2024) claims that the BPX morphology also depend on the bar angular momentum.
@@ 2 Simulation of N-body galaxy model
We generate the Milky Way-like galaxy N-body model using GAlIC (Yurin & Springel, 2014) code. The galaxy model has a stellar disk and dark matter halo components with a total galaxy mass of 6.38 $\times$ $10^{11}M_{\sun}$. There are 1 million particles in each component of the galaxy model. The stellar disk mass fraction is 0.1 of the total galaxy mass. The galaxy has an exponential disk with radial disk scale length $R_{d}=2.9$ kpc. The Hernquist dark matter halo, with an inner profile similar to the NFW halo (Springel et al., 2005), has a virial radius of $140$ kpc. The galaxy model used in this work is taken from our earlier work (Kataria & Shen, 2022) which is referred to as the S075 model in the previous work. We refer the reader to previous work for more details on the initial condition of the galaxy models used here.
@@ 3 Results
The top panel of Figure 1 shows the time evolution of bar strength. The bar strength continuously increases with time during the initial growth of the bar until 1.43 Gyr. The bar strength shows an abrupt decrease until 1.45 Gyr and it starts increasing continuously again until 2.2 Gyr. The bar strength does not grow between 2.2 to 3.2 Gyr though shows a short plateau distribution in this interval. Then beyond 3.2 Gyr bar strength grows and saturates in the secular evolution of the disk. Figure 2 shows the growth of bar length with time. The bar length decreases between 1.37 to 1.54 Gyr and shows a plateau between 2.2 to 3.2 Gyr like the bar strength evolution with time. The bar length continuously increases with time in secular evolution, unlike bar strength which saturates. We show the face-on and edge-on maps of bar evolution throughout the simulations in Figure 13 and discuss in Appendix A. The second panel of Figure 1 shows the time evolution of the pattern speed of the bar after the bar has formed ($A_{2}/A_{0}>0.35$). We find that pattern speed decreases with time after the bar gains strength by transferring angular momentum to the outer disk and dark matter halo (Kataria & Shen, 2022). The pattern speed shows a sharp decrease at 1.45 Gyr coinciding with the time of the sudden decrease in bar strength. In the secular phase of bar evolution, pattern speed decreases smoothly until the end of the simulation. The third panel of Figure 1 shows the time evolution of the vertical buckling strength which shows two clear peaks at 1.45 Gyr and 2.68 Gyr followed by small oscillations in secular evolution. These two buckling events, differing in 30 % in their strength, are represented by dashed and dotted vertical lines respectively which have the above-mentioned imprints on the bar strength and pattern speed evolution. The last panel of Figure 1 shows the time evolution of $\sigma_{z}/\sigma_{r}$ which is calculated within $2R_{d}$ region of the disk. The $\sigma_{z}/\sigma_{r}$ shows a decrease as the bar grows in the disk and results in the increased radial dispersion of the disk. Afterwards, there is an abrupt and huge increase in the $\sigma_{z}/\sigma_{r}$ value ($\approx$ 0.65 to 0.77) during the first buckling at 1.45 Gyr. Further, $\sigma_{z}/\sigma_{r}$ starts decreasing until the second buckling. In the second buckling event at 2.68 Gyr, there is a slow and relatively smaller increase in $\sigma_{z}/\sigma_{r}$ values ($\approx$ 0.68 to 0.72). In the secular phase of bar evolution, $\sigma_{z}/\sigma_{r}$ decreases and saturates at the end of the simulation.
@@ 3.1 Buckling events, corresponding BPX shape lengths and kinematic signatures
Similarly, we examine the face-on stellar density with isodensity contours evolution and the quadrupolar patterns in the mean $V_{z}$ maps evolution during the second buckling in Figures 7 and 8 respectively. We find that the semi-major axis of the bar remains the same ($\approx$ 6.2 kpc) during the whole secondary buckling event. Isodensity contours in Figure 7 show the weak spiral modes in the outer region of the disk which are not connected to the edge of the bar ($\approx$ 10 kpc). The quadrupolar patterns of the mean $V_{z}$ show a weak signal at the beginning of the buckling which becomes strengthened during the peak of buckling and fades away at around 3.08 Gyr. The region encompassing the quadrupolar patterns increases with radius from the beginning to the end of buckling event though remains confined within the semi-major axis of bar. In comparison to the first buckling, these quadrupolar patterns during the second buckling are extended, non-symmetric, and less pronounced compared to mean $V_{z}$ of the disk background.
@@ 3.2 Time Evolution of Fourier modes ($m=$ 2 and 4) throughout the disk
We observe that two buckling events have pronounced effects on the BPX shapes of bars as seen in Figure 4. In order to understand the impact of buckling throughout the disk we plot the $m=$ 2 and 4 modes in concentric rings of width equal to 0.75 kpc throughout the disk in Figures 9 and 10 respectively. The left panel of Figure 9 shows that the first buckling does not affect the growth of $A_{2}/A_{0}$ mode in the inner disk region (radius $<$ 6 kpc) while the second buckling shows a sudden dip in $A_{2}/A_{0}$ mode in this region. The division of the inner and outer regions is motivated by the difference in the time evolution of the overall pattern of $A_{2}/A_{0}$ and the dividing radius coincides with the typical bar semi-major axis length during the second buckling event. The right panel of Figure 9 shows that two buckling events have reverse trend, for a sudden dip in the time evolution of $A_{2}/A_{0}$ mode, in the outer disk (radius $>$ 6 kpc) compared to the inner part of the disk. This is interesting to observe that the first buckling which is confined to the central region of the disk (radius $<$ 3 kpc) affects the $m=2$ mode at the outer region of the disk. This is correlated with the growth and decay of spiral arms at the edge of bar (bar semi-major axis $\approx$ 4 kpc) during the first buckling as seen in Figure 5. On the other hand, the second buckling which spreads across the larger regions of the disk (radius $\approx$ 5 kpc) affects the $m=$ 2 mode of the central region of the disk. This is correlated with the inner region of the bar showing peanut features during the end of this buckling event as seen in Figure 7 at 3.08 Gyr. The time span of buckling events correlates inversely with the length scale of the disk where $m=$ 2 mode shows a sudden dip during its time evolution. For example the shorter time span of the first buckling shows a dip in the $m=$ 2 mode at large scale lengths of the disk and vice-versa. These modes show particular isophotal features in the face-on density maps which depend on the sequence or time span of buckling.
@@
Figure 9: Left panel shows the time evolution of m=2 Fourier mode ($A_{2}/A_{0}$) with time in concentric rings of increasing radius until 6 kpc of disk radius. The right panel show the same m=2 mode for concentric rings at a radius ranging from 6 kpc to 15 kpc. Here vertical dashed line corresponds to the first buckling ($\approx$ 1.45 Gyr) and the vertical dotted line corresponds to the second buckling event ($\approx$ 2.68 Gyr).
@@ 4 Discussion
Observations show that only a handful of galaxies (only 8 galaxies) have been caught in ongoing buckling action (Erwin & Debattista, 2016; Li et al., 2017; Xiang et al., 2021). Erwin & Debattista (2016) report two ongoing buckling events in their sample of 44 galaxies using photometric and kinematic signatures which suggest the buckling fraction to be around 4.5 %. Their toy model, which correctly predicts the frequency of buckling events with their observations, assumes the buckling timescale to be wide-ranged between 0.5 to 1 Gyr. Further, this study is limited by the sample size for a robust accounting of the frequency of ongoing buckling and bound on the timescale of buckling events. In a recent advancement that involves a novel approach, Xiang et al. (2021) report 5 ongoing buckling in a sample of 434 galaxies by detecting quadrupolar patterns in line-of-sight (LOS) velocities in the MaNGA survey having integral field spectroscopic data (IFU). This study also conveys the message of missing detection of buckling events in their sample due to limits on signal-to-noise ratios. Their toy model, based on a semi-analytical bar evolution model, predicts a lower bound on the buckling timescale to be 130 Myr. These semi-analytical calculations do not account for the secondary buckling events which can affect the detection probability and hence the lower bound on the buckling timescales. In our N-body model, the bar undergoes two buckling episodes lasting for around 220 Myr and 1 Gyr over a time span of 9 Gyr of bar evolution. Therefore, the detection probability of buckling events must be higher than the recently observed probability of these events, given the bar can last for buckling around an order of magnitude larger than predicted by Xiang et al. (2021).
@@
The time span of buckling and a number of successive buckling events are the key parameters for spotting the ongoing buckling action of bars in observed galaxies. In this work, we have looked into the role of successive buckling events on the potential observable properties of bar and disk. The kinematic observable signature involves the presence of quadrupolar patterns on the LOS velocity in the near face-on projection of disk. The longevity of these signatures is directly proportional to the time of buckling as evident from Figures 6 and 8. Further, the shorter time span of buckling affects $m=$ 2 and $m=$ 4 Fourier modes at larger scale lengths of disk. The observable signatures involve the growth and decay of spiral arms at the edge of bar ($m$=2) as well as the pronounced box shape of disk ($m$=4) during the first buckling event as seen in Figure 5. This can provide secondary signatures on the time spans of the buckling or successive buckling events in the photometric observations.
@@
Collier (2020) discussed the role of violent buckling compared to non-violent one where violent buckling has a lower timescale compared to non-violent one. The violent buckling does not destroy the bar and washes out quickly without affecting the bar in secular evolution while the non-violent one destroys the bar by sustaining bending modes for longer times. The current work has only violent buckling events ($A_{Buckle}>0.05$) given that do not destroy the bar irrespective of the time span of the bending mode. This highlights that the buckling mode lasting longer in the bar is not sufficient to destroy the bar.